OpenStudy (anonymous):

Simplify this trigonometry expression. (1-sinø) (1+sinø) csc^2ø + 1

OpenStudy (anonymous):

I got sec ^2ø for the answer, I don't know if it's correct.

OpenStudy (tkhunny):

Have you noticed that \((1-\sin(\theta))(1+\sin(\theta))\) is a sum and a difference that can be multiplied to a difference of squares?

OpenStudy (anonymous):

is csc^2ø correct?